# Sample-Efficient Constrained Reinforcement Learning with General Parameterization

### v1

@@ 4.3 Global Convergence of the Lagrange
Combining Lemma 3, 4 and using the expected gradient errors provided by Lemma 6, we bound the average Lagrange optimality gap as a function of tunable parameters $H$, $K$ and $\zeta$.
@@
Consider the same setup and the choice of parameters described in Lemma 3-- 6. The following inequality holds if assumptions 2--4 are met.
$\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{(}J_{\mathrm{L},\rho}(\pi^{*},\lambda_{k})-J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\bigg{)}\leq\sqrt{\epsilon_{\mathrm{bias}}}+\left[\dfrac{G\sqrt{C}(1+\lambda_{\max})}{\sqrt{\mu_{F}}(1-\gamma)^{2}}\right]\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)\\ &+B\eta\left(1+\dfrac{G^{4}}{\mu_{F}^{2}}\right)\left[\dfrac{22\sigma^{2}\mathrm{d}}{\mu_{F}H}+\exp\left(-\dfrac{\mu_{F}}{20G^{2}}H\right)\left[\dfrac{C(1+\lambda_{\max})^{2}}{\mu_{F}(1-\gamma)^{4}}\right]\right]+\left(\dfrac{4BG^{2}(1+\lambda_{\max})}{\mu_{F}^{2}(1-\gamma)}\right)\dfrac{1}{K}\\ &+\left(\dfrac{2BG^{2}}{\mu_{F}^{2}(1-\gamma)^{2}}\right)\zeta+\frac{1}{\eta K}\mathbf{E}_{s\sim d^{\pi^{*}}_{\rho}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]\end{split}$ (30)
@@ 4.4 Decoupling the Objective and the Constraint Violation Rates
Consider the same setup and the choice of parameters given in Lemma 3-- 6. Assume $\zeta=\lambda_{\max}(1-\gamma)/\sqrt{K}$ where $\zeta$ denotes the learning rate of the Lagrange update (21), and $\lambda_{\max}=2/[(1-\gamma)c_{\mathrm{slater}}]$. For arbitrary $\epsilon>0$, the following inequalities hold
@@
whenever assumptions 1--4 are met, $H=\mathcal{O}((1-\gamma)^{-3}\epsilon^{-1})$ and $K=\mathcal{O}((1-\gamma)^{-4}\epsilon^{-2})$. Therefore, the sample complexity to ensure (31) is $\mathcal{O}((1-\gamma)^{-1}HK)=\mathcal{O}((1-\gamma)^{-8}\epsilon^{-3})$.
Theorem 1 dictates that, with appropriate choice of the parameters, the rate of convergence of the objective and that of constraint violation can be bounded as $\mathcal{O}(\sqrt{\epsilon_{\mathrm{bias}}}+\epsilon)$ with $\mathcal{O}((1-\gamma)^{-8}\epsilon^{-3})$ samples. This beats the SOTA $\tilde{\mathcal{O}}(\epsilon^{-4})$ sample complexity of [4].
@@
Note the importance of the nested expectation in the first-order term $\mathbf{E}\|\mathbf{E}[\omega_{k}|\theta_{k},\lambda_{k}]-\omega_{k}^{*}\|$ in Lemma 3. Had this term been $\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|$, it would have resulted in an average Lagrange optimality gap of $\mathcal{O}\left(H^{-0.5}+K^{-1}+\zeta\right)$ in Corollary 1 which would have lead to a sample complexity of $\mathcal{O}(\epsilon^{-4})$. However, due to the improved first-order term obtained via our careful analysis, the gap improves to $\mathcal{O}\left(H^{-1}+K^{-1}+\zeta\right)$ resulting in $\mathcal{O}(\epsilon^{-3})$ sample complexity.
Remark 4.
Note that the RHS of (30) is a decreasing function of $\zeta$, which, unlike $H$ and $K$, does not contribute towards the sample complexity. One, therefore, might want to take $\zeta$ as small as possible. However, the proof of Theorem 1 shows that if $\zeta$ is taken too small, then the constraint violation cannot be contained by $\epsilon$. The optimal middle ground is achieved by taking $\zeta=\mathcal{O}(K^{-0.5})$.
@@ 5 Conclusions and Limitations
This paper considers the problem of learning a CMDP where the goal is to maximize the objective value function while guaranteeing that the cost value exceeds a predefined threshold. We propose an acceleration-based primal-dual natural policy gradient algorithm that ensures $\epsilon$ optimality gap and $\epsilon$ constraint violation with $\mathcal{O}(\epsilon^{-3})$ sample complexity. This improves upon the previous state-of-the-art sample complexity of $\mathcal{O}(\epsilon^{-4})$. The main limitation of our work is that the obtained sample complexity is still larger than the theoretical lower bound. Future works include improving the sample complexity to the optimal value of $\mathcal{O}(\epsilon^{-2})$, extending the idea of acceleration to nonlinear reinforcement learning problems, designing parameter-free CMDP algorithms, etc.
@@ Appendix D Proof of Lemma 4
Recall from Lemma 2(b) that $J_{\mathrm{L},\rho}(\cdot,\lambda)$ is $L(1+\lambda)$-smooth. Therefore,
$\begin{split}&J_{\mathrm{L},\rho}(\theta_{k+1},\lambda_{k})\geq J_{\mathrm{L}}(\theta_{k},\lambda_{k})+\left<\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k}),\theta_{k+1}-\theta_{k}\right>-\frac{L(1+\lambda_{k})}{2}\|\theta_{k+1}-\theta_{k}\|^{2}\\ &\overset{(a)}{=}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})+\eta\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})^{T}\omega_{k}-\frac{L(1+\lambda_{k})\eta^{2}}{2}\|\omega_{k}\|^{2}\\ &=J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})+\eta\langle\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k}),\omega_{k}^{*}\rangle+\eta\langle\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k}),\omega_{k}-\omega_{k}^{*}\rangle\\ &\hskip 65.44142pt-\frac{L(1+\lambda_{k})\eta^{2}}{2}\|\omega_{k}-\omega_{k}^{*}+\omega_{k}^{*}\|^{2}\\ &\overset{}{\geq}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})+\eta\langle\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k}),F_{\rho}(\theta_{k})^{\dagger}\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\rangle+\eta\langle\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k}),\omega_{k}-\omega_{k}^{*}\rangle\\ &\hskip 65.44142pt-L(1+\lambda_{k})\eta^{2}\|\omega_{k}-\omega_{k}^{*}\|^{2}-L(1+\lambda_{k})\eta^{2}\|\omega_{k}^{*}\|^{2}\\ &\overset{(b)}{\geq}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})+\dfrac{\eta}{G^{2}}\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}+\eta\langle\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k}),\omega_{k}-\omega_{k}^{*}\rangle\\ &\hskip 65.44142pt-L(1+\lambda_{k})\eta^{2}\|\omega_{k}-\omega_{k}^{*}\|^{2}-L(1+\lambda_{k})\eta^{2}\|F_{\rho}(\theta_{k})^{\dagger}\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}\\ &\overset{(c)}{\geq}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})+\left(\dfrac{\eta}{2G^{2}}-\dfrac{L(1+\lambda_{k})\eta^{2}}{\mu_{F}^{2}}\right)\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}\\ &+\dfrac{\eta}{2G^{2}}\underbrace{\left[\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}+2G^{2}\langle\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k}),\omega_{k}-\omega_{k}^{*}\rangle+G^{4}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right]}_{=\left\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})+G^{2}(\omega_{k}-\omega_{k}^{*})\right\|^{2}\geq 0}\\ &-\left(L(1+\lambda_{k})\eta^{2}+\dfrac{\eta G^{2}}{2}\right)\|\omega_{k}-\omega_{k}^{*}\|^{2}\\ &\geq J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})+\left(\dfrac{\eta}{2G^{2}}-\dfrac{L(1+\lambda_{k})\eta^{2}}{\mu_{F}^{2}}\right)\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}\\ &\hskip 65.44142pt-\left(L(1+\lambda_{k})\eta^{2}+\dfrac{\eta G^{2}}{2}\right)\|\omega_{k}-\omega_{k}^{*}\|^{2}\end{split}$ (42)
where (a) follows from the update equation $\theta_{k+1}=\theta_{k}+\eta\omega_{k}$ and (b), (c) utilize Assumption 2 and 4 respectively. Adding $J_{\mathrm{L},\rho}(\theta_{k+1},\lambda_{k+1})$ on both sides of the above inequality, we obtain the following.
$\begin{split}&J_{\mathrm{L},\rho}(\theta_{k+1},\lambda_{k+1})-J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\geq J_{\mathrm{L},\rho}(\theta_{k+1},\lambda_{k+1})-J_{\mathrm{L},\rho}(\theta_{k+1},\lambda_{k})\\ &+\left(\frac{\eta}{2G^{2}}-\dfrac{L(1+\lambda_{k})\eta^{2}}{\mu_{F}^{2}}\right)\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}-\left(\frac{\eta G^{2}}{2}+L(1+\lambda_{k})\eta^{2}\right)\|\omega_{k}-\omega^{*}_{k}\|^{2}\end{split}$ (43)
Observe that,
$\begin{split}J_{\mathrm{L},\rho}(\theta_{k+1},\lambda_{k+1})-J_{\mathrm{L},\rho}(\theta_{k+1},\lambda_{k})&=(\lambda_{k+1}-\lambda_{k})J_{c,\rho}(\theta_{k+1})\\ &\overset{(a)}{\geq}-\dfrac{1}{1-\gamma}|\lambda_{k+1}-\lambda_{k}|\overset{(b)}{\geq}-\dfrac{\zeta}{1-\gamma}\left|\hat{J}_{c,\rho}(\theta_{k})\right|\end{split}$ (44)
where (a) utilizes the fact that $|J_{c,\rho}(\theta_{k+1})|\leq 1/(1-\gamma)$ whereas (b) results from (21) and contractive property of the projection operator $\mathcal{P}_{\Lambda}$. Combining (43) and (44), and using $\lambda_{k}\leq\lambda_{\max}$, one obtains,
$\begin{split}&J_{\mathrm{L},\rho}(\theta_{k+1},\lambda_{k+1})-J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\geq-\dfrac{\zeta}{1-\gamma}\left|\hat{J}_{c,\rho}(\theta_{k})\right|\\ &+\left(\frac{\eta}{2G^{2}}-\dfrac{L(1+\lambda_{\max})}{\mu_{F}^{2}}\eta^{2}\right)\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}-\left(\frac{\eta G^{2}}{2}+L(1+\lambda_{\max})\eta^{2}\right)\|\omega_{k}-\omega^{*}_{k}\|^{2}\end{split}$ (45)
Choosing $\eta=\frac{\mu_{F}^{2}}{4G^{2}L(1+\lambda_{\max})}$, and taking a sum over $k\in\{0,\cdots,K-1\}$, we arrive at,
$\begin{split}&\dfrac{\mu_{F}^{2}}{16G^{4}L(1+\lambda_{\max})}\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}\right)\leq\dfrac{\zeta}{1-\gamma}\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\left|\hat{J}_{c,\rho}(\theta_{k})\right|\right)\\ &+\dfrac{1}{L(1+\lambda_{\max})}\left(\dfrac{\mu_{F}^{2}}{8}+\dfrac{\mu_{F}^{4}}{16G^{4}}\right)\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)+\dfrac{J_{\mathrm{L},\rho}(\theta_{K},\lambda_{K})-J_{\mathrm{L},\rho}(\theta_{0},\lambda_{0})}{K}\end{split}$ (46)
Note that $|J_{\mathrm{L},\rho}(\theta_{K},\lambda_{K})-J_{\mathrm{L},\rho}(\theta_{0},\lambda_{0})|\leq 2(1+\lambda_{\max})/(1-\gamma)$. Moreover, following the sampling procedure described in Algorithm 19, we deduce, $\mathbf{E}|\hat{J}_{c,\rho}(\theta_{k})|\leq\sum_{t=0}^{\infty}(1-\gamma)(t+1)\gamma^{t}=1/(1-\gamma)$. This leads us to the following.
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\nabla_{\theta}J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\|^{2}&\leq(2G^{4}+\mu_{F}^{2})\left(\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|^{2}\right)\\ &+\left(\dfrac{32G^{4}L(1+\lambda_{\max})^{2}}{\mu_{F}^{2}(1-\gamma)}\right)\dfrac{1}{K}+\left(\dfrac{16G^{4}L(1+\lambda_{\max})}{\mu_{F}^{2}(1-\gamma)^{2}}\right)\zeta\end{split}$ (47)
This completes the proof. ∎
Appendix E Proof of Lemma 5
Proof.
@@ Appendix E Proof of Lemma 5
$\begin{split}&\mathbf{E}\left[\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega_{\theta,\lambda}^{*},\theta,\lambda)\otimes\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega_{\theta,\lambda}^{*},\theta,\lambda)\right]\\ &=\mathbf{E}_{(s,a)\sim\nu_{\rho}^{\pi^{*}}}\bigg{[}\mathbf{E}\bigg{[}\underbrace{\nabla_{\theta}\log\pi_{\theta}(a|s)\cdot\omega^{*}_{\theta,\lambda}-\dfrac{1}{1-\gamma}\hat{A}_{\mathrm{L},\lambda}^{\pi_{\theta}}(s,a)}_{\triangleq\zeta_{\theta,\lambda}(s,a)}\bigg{]}^{2}\nabla_{\theta}\log\pi_{\theta}(a|s)\otimes\nabla_{\theta}\log\pi_{\theta}(a|s)\bigg{]}\end{split}$
@@
Appendix F Proof of Lemma 6
@@
Appendix G Proof of Theorem 1
Applying the inequalities $\exp(-x)<\frac{1}{x}$, $\forall x>0$, $B/(1-\gamma)^{2}\leq L\leq(B+2G^{2})/(1-\gamma)^{3}$, and substituting the value of $\eta$ stated in Lemma 4, we can rewrite (30) as follows.
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{(}J_{\mathrm{L},\rho}(\pi^{*},\lambda_{k})&-J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\bigg{)}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{H}+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{3}}\dfrac{1}{K}+\dfrac{f_{2}}{(1-\gamma)^{2}}\zeta\end{split}$ (61)
The terms $f_{0},f_{1},f_{2}$ are defined below.
$\begin{split}&f_{0}\triangleq\dfrac{40G^{3}\sqrt{C}}{\mu_{F}^{3/2}}+\left(\dfrac{\mu_{F}^{2}}{G^{2}}+G^{2}\right)\left[\dfrac{11\mathrm{d}}{\mu_{F}}\left(\dfrac{G^{4}}{\mu_{F}^{2}}+16\right)+\dfrac{5G^{2}C}{\mu_{F}^{2}}\right],\\ &f_{1}\triangleq\dfrac{4G^{2}}{\mu_{F}^{2}}\left[B+(B+2G^{2})\mathbf{E}_{s\sim d^{\pi^{*}}_{\rho}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]\right],\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ f_{2}\triangleq\dfrac{2BG^{2}}{\mu_{F}^{2}}\end{split}$ (62)

### v3

@@ 4.3 Global Convergence of the Lagrange
Combining Lemma 3, (23) and using the expected gradient errors provided by Lemma 5, we bound the average Lagrange optimality gap as a function of tunable parameters $H$ and $K$ as stated in the following corollary.
@@
Consider the same setup and the choice of parameters described in Lemma 3-- 5. The following inequality holds if assumptions 2--4 are met.
$\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{(}J_{\mathrm{L},\rho}(\pi^{*},\lambda_{k})-J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\bigg{)}\leq\sqrt{\epsilon_{\mathrm{bias}}}+\left[\dfrac{G\sqrt{C}(1+\lambda_{\max})}{\sqrt{\mu_{F}}(1-\gamma)^{2}}\right]\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)\\ &+B\eta\left[\dfrac{22\sigma^{2}\mathrm{d}}{\mu_{F}H}+\exp\left(-\dfrac{\mu_{F}}{20G^{2}}H\right)\left[\dfrac{C(1+\lambda_{\max})^{2}}{\mu_{F}(1-\gamma)^{4}}\right]+\dfrac{G^{2}(1+\lambda_{\max})^{2}}{\mu_{F}^{2}(1-\gamma)^{4}}\right]\\ &+\frac{1}{\eta K}\mathbf{E}_{s\sim d^{\pi^{*}}_{\rho}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]\end{split}$ (29)
Corollary 1 bounds the optimality error of the Lagrange function as $\mathcal{O}(\sqrt{\epsilon_{\mathrm{bias}}}+\exp\left(-C_{0}H\right)+\eta+\frac{1}{\eta K})$ where $C_{0}$ is some problem specific constant. Interestingly, the dual learning parameter, $\zeta$ does not appear in this bound. However, the next section shows that $\zeta$ plays a pivotal role in deciding the objective and constraint violation rates.
@@ 4.4 Decoupling the Objective and the Constraint Violation Rates
Consider the same setup and the choice of parameters given in Lemma 3-- 5. Assume $\eta=(1-\gamma)^{2}(1+\lambda_{\max})^{-1}/\sqrt{K}$, $\zeta=\lambda_{\max}(1-\gamma)/\sqrt{K}$, and $\lambda_{\max}=2/[(1-\gamma)c_{\mathrm{slater}}]$. For sufficiently small $\epsilon>0$, the following inequalities hold
@@
whenever assumptions 1--4 are met, $H=\mathcal{O}(\log(\epsilon^{-1}))$ and $K=\mathcal{O}((1-\gamma)^{-6}\epsilon^{-2})$. Therefore, the sample complexity to ensure (30) is $\mathcal{O}((1-\gamma)^{-1}HK)=\tilde{\mathcal{O}}((1-\gamma)^{-7}\epsilon^{-2})$. It is to be clarified that the $(1-\gamma)^{-1}$ factor in the sample complexity calculation appears due to the fact that it requires $\mathcal{O}((1-\gamma)^{-1})$ samples on an average to obtain a gradient estimate via Algorithm 19.
Theorem 1 dictates that, with appropriate choice of the parameters, the rate of convergence of the objective and that of constraint violation can be bounded as $\mathcal{O}(\sqrt{\epsilon_{\mathrm{bias}}}+\epsilon)$ with $\tilde{\mathcal{O}}((1-\gamma)^{-7}\epsilon^{-2})$ samples. This beats the SOTA $\tilde{\mathcal{O}}(\epsilon^{-4})$ sample complexity of [4] and achieves the theoretical lower bound. Our derived sample complexity is also dependent on $c_{\mathrm{slater}}$. However, this is not explicitly mentioned in Theorem 1. Interested readers can find such details in the appendix.
@@
Note the importance of the nested expectation in the first-order term $\mathbf{E}\|\mathbf{E}[\omega_{k}|\theta_{k},\lambda_{k}]-\omega_{k}^{*}\|$ in Lemma 3. Lemma 5 bounds this term as $\mathcal{O}(\exp(-C_{0}H))$ where $C_{0}$ is a problem dependent constant. Other terms in the Lagrange optimality bound (Lemma 3) can be expressed as $\mathcal{O}(\eta+\frac{1}{\eta K})$. Moreover, following the proof of Theorem 1, one sees that decoupling the objective optimality error and constraint violation bounds incurs additional $\mathcal{O}(\zeta+\frac{1}{\zeta K})$ terms. Choosing $\eta,\zeta$ as prescribed in Theorem 1, makes both the objective and constraint violation errors as $\mathcal{O}(\sqrt{\epsilon_{\mathrm{bias}}}+\exp(-C_{0}H)+K^{-0.5})$. This allows us to take $H=\tilde{\mathcal{O}}(1)$ and $K=\mathcal{O}(\epsilon^{-2})$, leading to $\tilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity. Had the first-order term been $\mathbf{E}\|\omega_{k}-\omega_{k}^{*}\|$, it would have resulted in $\mathcal{O}(\sqrt{\epsilon_{\mathrm{bias}}}+H^{-0.5}+K^{-0.5})$ objective and constraint violation errors which would have lead to $\mathcal{O}(\epsilon^{-4})$ sample complexity.
@@ 5 Conclusions and Limitations
This paper considers the problem of learning a CMDP where the goal is to maximize the objective value function while guaranteeing that the cost value exceeds a predefined threshold. We propose an acceleration-based primal-dual natural policy gradient algorithm that ensures $\epsilon$ optimality gap and $\epsilon$ constraint violation with $\tilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity. This improves upon the previous state-of-the-art sample complexity of $\mathcal{O}(\epsilon^{-4})$ and achieves the theoretical lower bound. Future works include applying the idea of acceleration-based NPG to improve sample complexities in other related domains of constrained reinforcement learning, e.g., non-linear CMDP, average reward CMDP, etc.
@@ Appendix D Proof of Lemma 4
$\begin{split}&\mathbf{E}\left[\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega_{\theta,\lambda}^{*},\theta,\lambda)\otimes\hat{\nabla}_{\omega}L_{\nu^{\pi_{\theta}}_{\rho}}(\omega_{\theta,\lambda}^{*},\theta,\lambda)\right]\\ &=\mathbf{E}_{(s,a)\sim\nu_{\rho}^{\pi_{\theta}}}\bigg{[}\mathbf{E}\bigg{[}\underbrace{\nabla_{\theta}\log\pi_{\theta}(a|s)\cdot\omega^{*}_{\theta,\lambda}-\dfrac{1}{1-\gamma}\hat{A}_{\mathrm{L},\lambda}^{\pi_{\theta}}(s,a)}_{\triangleq\zeta_{\theta,\lambda}(s,a)}\bigg{]}^{2}\nabla_{\theta}\log\pi_{\theta}(a|s)\otimes\nabla_{\theta}\log\pi_{\theta}(a|s)\bigg{]}\end{split}$
@@
Appendix E Proof of Lemma 5
@@
Appendix F Proof of Theorem 1
Applying the inequalities $H\geq 1$, $\exp(-(\mu_{F}/20G^{2})H)\leq 1$ and substituting the values of $\sigma^{2}$ and $\eta$ (as stated in Lemma 4 and Theorem 1 respectively), we can rewrite (29) as follows.
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{(}J_{\mathrm{L},\rho}(\pi^{*},\lambda_{k})&-J_{\mathrm{L},\rho}(\theta_{k},\lambda_{k})\bigg{)}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}\end{split}$ (54)
The terms $f_{0}$ and $f_{1}$ are defined below.
$\begin{split}&f_{0}\triangleq\dfrac{G\sqrt{C}}{\sqrt{\mu_{F}}},\\ &f_{1}\triangleq B\left[\dfrac{44\mathrm{d}}{\mu_{F}}\left(\dfrac{G^{4}}{\mu_{F}^{2}}+16\right)+\dfrac{C}{\mu_{F}}+\dfrac{G^{2}}{\mu_{F}^{2}}\right]+\mathbf{E}_{s\sim d^{\pi^{*}}_{\rho}}[KL(\pi^{*}(\cdot|s)\|\pi_{\theta_{0}}(\cdot|s))]\end{split}$ (55)